OpenStudy (anonymous):

Solve x - 6(3 - x) = 7(3 + x)

OpenStudy (saifoo.khan):

x - 18 +6x = 21 + 7x

OpenStudy (anonymous):

Wow, that is the final answer?

OpenStudy (anonymous):

x=39+x

OpenStudy (saifoo.khan):

no, that isnt.

OpenStudy (saifoo.khan):

x - 18 +6x = 21 + 7x 7x -18 = 21 + 7x No solution!

OpenStudy (anonymous):

ok, so it is no solution or -39 = 0

OpenStudy (saifoo.khan):

Yea!

OpenStudy (anonymous):

x-6(3-x)=7(3-x) x-18+6x=21-7x 7x-18=21-7x no answer because x is cancled